QL=-L^3+8L^2+12 then find (a)optimal level of labour employment (b)optimal out put level(c) labour employment at which maximum out put is produced (hint;TPL is maximum at MPL=0)
1
Expert's answer
2018-02-10T06:19:07-0500
QL=-L^3+8L^2+12 then find
(a) optimal level of labour employment is at:
QL' = 0,
-3L^2 + 16L = 0,
L(16 - 3L) = 0,
L = 0 or L = 5.33.
(b) optimal output level can't be found without additional data.
(c) labour employment at which maximum out put is produced can't be found without additional data.